OpenStudy (anonymous):

Does anybody know how to use Green's Theorem to evaluate the line integral.......

OpenStudy (anonymous):

\[\int\limits_{?}^{?}xy dx + y dy \] where C is the unit circle.

OpenStudy (anonymous):

Do you know the statement of green's theorem?

OpenStudy (anonymous):

yes i found P=xy and Q=y with dp/dy=x and dq/dx=0 so dq/dy - dp/dy =-x

OpenStudy (anonymous):

so I got the double integral to be \[\int\limits_{0}^{2\pi}\int\limits_{0}^{1}-\cos \theta r drd \theta \] after i converted to polar coordinates

OpenStudy (anonymous):

when i go through the integration i get 0 which i don't think is correct

OpenStudy (anonymous):

Shouldn't x=rcos(theta)

OpenStudy (anonymous):

yes i just wrote it as \[\cos \theta r dr d \theta \] and since it is a -x it comes out to \[-rcos \theta dr d \theta \]

OpenStudy (anonymous):

There should also be an r from the change of variables I believe. Anyway have you tried evaluating the line integral normally to check. Unit circle can be parametrized easily.

OpenStudy (anonymous):

i thought since the parametrization of the unit circle is \[x= \cos \theta\] and \[y=\sin \theta \] with the r in front = to 1 since it is the unit circle you just have to include 1 r for the change of variables. I haven't tried evaluating the line integral normally. Not sure I remember how....lol

OpenStudy (anonymous):

You arn't parametrizing when doing an integral over D. You should do that when actually evaluating the line integral. You need rcos(theta) because you are varying r from 0 to 1 to cover the unit disk.

OpenStudy (anonymous):

ok i'm confused. how do you set up the integral using Green's Theorem?

OpenStudy (anonymous):

Just like you did. It's \[-\int_{R} x\ dydx\] where R is the unit disk. So you can forget that you are doing a line integral and just evaluate the double integral as you normally would. \[(x,y)=(r\cos(\theta), r\sin(\theta)); dydx=rdrd\theta\]

OpenStudy (anonymous):

so in this case what would the limits of integration be? \[-\int\limits_{0}^{2\pi}\int\limits_{0}^{1}x dydx\] ??

OpenStudy (anonymous):

Just like you have it. Except do the change of variables correctly.

OpenStudy (anonymous):

so it should be \[-\int\limits\limits_{0}^{2\pi}\int\limits\limits_{0}^{1} r ^{2}\cos \theta drd \theta \] ???

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

when i evaluate the integral i still get 0 which i don't think is correct

OpenStudy (anonymous):

Why isn't it correct?

OpenStudy (anonymous):

what does the integral represent? area? volume? and if so how can an area or volume be 0?

OpenStudy (anonymous):

Remember that integrals evaluated signed area and volumes. There are negative areas and volumes. You can check your answer by evaluating the original line integral.

OpenStudy (anonymous):

Ok thanks, I will need to evaluate the line integral later, I have to run some errands. Thanks for you help.